# envname-P envname#1

### v2

@@ 4.4. Preference-aware POI Retrieval
Ideally, we want the queried POIs best fit the positive subrequest while avoiding the negative subrequest. To achieve this, we first use the calculated positive embedding $\small\mathbf{e}^{\operatorname{pos}}_{i}$ to retrieve $\small k$ POIs $\small\mathcal{P}^{\operatorname{pos}}_{i}=\{p^{\operatorname{pos}}_{i,j}\}_{j=1}^{k}$ with top similarity scores $\small\mathcal{S}^{\operatorname{pos}}_{i}=\{s^{\operatorname{pos}}_{i,j}\}_{j=1}^{k}$ and embedding $\small\mathcal{E}^{\operatorname{pos}}_{i}\in\mathbb{R}^{k\times d}$, where $\small s^{\operatorname{pos}}_{i,j}$ and $\small s^{\operatorname{pos}}_{i,j}$ represent the $\small j$th POI and score for $\small i$th positive sub-request, respectively. Then, to ensure avoiding negative POIs, we compute the similarity scores of $\small\mathcal{E}^{\operatorname{pos}}_{i}$ and $\small\mathbf{e}^{\operatorname{neg}}_{i}$ and rerank the retrieved POIs base on the gap between POI-wise positive and negative similarity scores.
Suppose we use $\small\mathcal{E}\in\mathbb{R}^{N\times d}$ to denote the POI embeddings pre-computed over POI attributes that stored in the user-owned embedding database, the above process could be formulate as:
[…]
@@
where the $\small\operatorname{score}(\cdot)$ function measures embedding similarities, and the superscript $\small k$ indicates that the function only returns the top-$\small k$ results based on the similarity scores. The $\small\operatorname{rank}(\cdot)$ reorders POIs based on their similarity scores from highest to lowest.
@@
Lastly, we select the POIs with the highest top-$\small k$ summed-scores from the unioned set of all retrieved POIs to form the final retrieved POI set $\small\mathcal{P}^{rt}$ for the user request $\small r$:
(7) $\small\mathcal{P}^{rt},\mathcal{S}^{rt}=\operatorname{rank}^{k}\left(\cup_{i=1}^{|\mathcal{R}|}\left(\mathcal{P}_{i},\mathcal{S}_{i}\right)\right)\cup\left(\mathcal{P}^{\operatorname{must}},\mathcal{S}^{\operatorname{must}}\right),$
@@
where the set $\small\mathcal{S}^{\operatorname{must}}$ is assigned large values to ensure must-see POIs are included in the final itinerary by subsequent modules. During the union process, the scores of the same POI under different subrequests are summed to form the final score of that POI.
It is worth noting that avoiding negative POIs in the retrieval stage requires scoring across $\small\mathcal{E}^{\operatorname{neg}}_{i}$ and the entire user-owned embedding database, which may hinder the response times. Therefore in actual deployment, we focus on reranking retrieved positive POIs base on negative and positive scores. Given that Eq. 7 ranks all retrieved POIs, our approach does not compromise retrieval effectiveness.
4.5. Cluster-aware Spatial Optimization
A spatially coherent itinerary could essentially enhance the travel experience, travelers tend to move from one cluster of POIs to another (Bolzoni et al., 2014). This behavior can be attributed to the efficiency and convenience of exploring POIs within a proximal area, reducing the time and effort involved in transit. In view of this, spatially filtering all retrieved POIs and sequencing their visitation order is essential in order to inform large language model which so far has exhibited limited capabilities for understanding layout and optimizing routes in multi-dimensional space. To achieve this, we initially compute spatial clusters of the retrieved POIs and then select candidate POIs based on geographical proximity and matching scores. We address cluster-aware spatial optimization by solving a hierarchical traveling salesman problem (Jiang et al., 2014).
4.5.1. Obtain Candidate POI via Spatial Clustering and Filtering
Given the set of retrieved POIs, $\small\mathcal{P}^{rt}$, we create a spatial proximity graph $\small G$ using a distance matrix $\small D$. In this graph, each node denotes a POI, and edges connect nodes (POIs) that are closer than a specific distance threshold $\small\tau$. A community detection algorithm is applied to $\small G$ so that all the nodes are divided into multiple mutually exclusiveclusters. Next, we iteratively select the clusters with the largest summed similarity scores derived from the PPR module until the total number of selected POIs reach a predefined threshold $\small N^{c}$, and the POIs in these clusters form the candidate POIs $\small\mathcal{P}^{c}$ for the user request $\small r$. The above process is detailed in Algo. 1.
Algorithm 1 Spatial Clustering and Selection of Candidate POIs
@@ 4.5.1. Obtain Candidate POI via Spatial Clustering and Filtering
1:Input: Retrieved POI set $\small\mathcal{P}^{rt}$ with scores $\small\mathcal{S}^{rt}$, Distance threshold $\small\tau$, Candidate POIs num threshold $\small N^{c}$
2:Output: Spatial Clusters $\small\mathcal{C}$, Candidate POIs $\small\mathcal{P}^{c}$
@@
5:for $\small p_{a}^{rt},p_{b}^{rt}\in V$ with $\small a\neq b$ do
6:  if $\small{distance}(p_{a}^{rt},p_{b}^{rt})<\tau$ then
7:   $\small E\leftarrow E\cup\{(p_{a}^{rt},p_{b}^{rt})\}$
@@
10:while $\small V\neq\emptyset$ do
11:  $\small c\leftarrow\text{largest clique in }G$
12:  $\small\mathcal{C}\leftarrow\mathcal{C}\cup\{c\}$; $\small V\leftarrow V\setminus c$
@@
15:for each cluster $\small c\in\mathcal{C}$ do
16:  $\small s^{rt}_{c}\leftarrow\sum_{p_{j}\in c}s_{j}^{rt}$
@@
18:Sort $\small\mathcal{C}$ in descending order of $\small\mathcal{S}^{c}=\{s^{rt}_{c}\}_{c=1}^{\mathcal{C}}$
19:while $\small|\mathcal{P}^{c}|<N^{c}$ do
20:  $\small c_{\text{max}}\leftarrow\arg\max_{c\in\mathcal{C}}s_{c}^{rt}$
21:  $\small\mathcal{P}^{c}\leftarrow\mathcal{P}^{c}\cup c_{\text{max}}$; $\small\mathcal{C}\leftarrow\mathcal{C}\setminus\{c_{\text{max}}\}$
@@
23:return $\small\mathcal{C},\mathcal{P}^{c}$
4.5.2. POI Ordering via Solving a Hierarchical TSP
After obtaining the spatial clusters $\small\mathcal{C}$, we further optimize the visitation order of the candidate POIs to ensure a spatially coherent itinerary. This process first involves determining the access order of each cluster. Then within each cluster, we determine the visitation order of POIs by solving TSPs with starting and ending POI constraints. Specifically, the start and end points for each cluster are selected based on the proximity of POIs in adjacent clusters as described in Fig. 2. We provide an Algo. 2 below that illustrates the procedure.
[…]
@@ 4.5.2. POI Ordering via Solving a Hierarchical TSP
1:Input: Spatial clusters $\small\mathcal{C}$, candidate POIs $\small\mathcal{P}^{c}$, distance matrix $\small D$
2:Output: Ordered list of candidate POIs $\small\mathcal{P}^{\text{order}}$
@@
5:for each cluster $\small c$ in $\small\mathcal{C}^{\text{order}}$ do
6:  $\small p_{\text{start}}^{c},p_{\text{end}}^{c}\leftarrow\text{GetClusterEndpoints}(c,\mathcal{C}^{\text{order}},D)$
7:  $\small c_{\text{path}}\leftarrow\text{SolveConstraintTSP}(c,p_{\text{start}}^{c},p_{\text{end}}^{c},D)$
8:  $\small\mathcal{P}^{\text{order}}\leftarrow\mathcal{P}^{\text{order}}\cup c_{\text{path}}$
@@
13:return $\small\mathcal{P}^{\text{order}}$
4.6. Itinerary Generation
@@
Selecting any subset from $\small\mathcal{P}^{\text{order}}$ guarantees a spatially coherent itinerary, but a high-quality itinerary should also comply with additional constraints, such as matching the user’s time availability and ensuring practicality. For example, it should avoid consecutive restaurant visits and ensure venues are visited at appropriate times, such as not planning bars in the morning or coffee shops late at night. Given the complex nature of these constraints, traditional optimization-based algorithms can become excessively complicated and lack variability (Yochum et al., 2020; Taylor et al., 2018), hindering system deployment and itinerary diversity. To address this, considering the advanced reasoning and planning capabilities of LLMs, we leverage LLMs to generate final itineraries that satisfy these diverse constraints.
Simplified prompt $\small\mathbb{P}_{IG}$ for final itinerary generation. See Appendix F.4 for full version. Please consider carefully and use the provided Candidate POIs list to craft a one-day itinerary in the form of an engaging and realistic travel story. ## Itinerary Information: - **Candidate POIs**: {p_order} 1: Time Cafe, Dongcheng District. Located in the bustling Nanluoguxiang, this cafe is a popular spot for photo enthusiasts and also a cat cafe. Surrounded by beautiful roses, its perfect for capturing memories.“ 2: - **Must-see POIs**: {r_must} - **Keyword Requirements**: {r_pos} - **Users Original Request**: {r} - **Start POI**: {p_start} - **End POI**: {p_end} ## Constraints: - Only choose POI from the **Candidate POI** list and in ascending order. - Bars should be at the end of the itinerary, while coffee shops should not be the last POI. ## Task Description: 1. You have approximately {hours} hours. Select suitable POIs from the **Candidate POIs** list in ascending order, ensuring your itinerary is selectively filtered and does not include all POIs from the list. 2. All POIs added to the itinerary must follow the ascending order of **Candidate POIs**. 3. Generate a JSON file containing all selected POIs. ## Output Format: {{ itinerary”: List of POIs, separated by ’->’” Overall reason”: Overall recommendation reason for the designed day trip itinerary”, pois”: {{ n”: Description and recommendation reason for each POI”, }} }}
@@
Specifically, the primary objective of this module is to effectively utilize LLM to select an optimal subset from $\small\mathcal{P}^{\text{order}}$, which closely aligns with user requests while adhering to various constraints. This process can be formally defined as follows:
[…]
Utilizing an LLM to generate the final travel itinerary allows for a balance among various considerations. Additionally, leveraging the understanding and reasoning capabilities of LLM, we can create a time-appropriate travel itinerary tailored to the user requests. Specifically, we leverage a simple time indication prompt $\small\mathbb{P}_{TI}$ to indicate the travel time (in hours) of the itinerary, and the results can be integrated into the “Task Description” part of the prompt $\small\mathbb{P}_{IG}$. We provide the complete $\small\mathbb{P}_{TI}$ and further details on the travel time considerations of the itinerary in Appendix F.2.
5. Experiments
To evaluate the performance of ItiNera for OUIP, we conduct extensive experiments to answer the following research questions: RQ1, why does directly using LLM to generate travel itineraries not guarantee a good user experience, and what advantages can ItiNera offer in comparison? RQ2, how do the various components of ItiNera enhance the quality of the generated itineraries? RQ3, how well does ItiNera serve the OUIP service at TuTu in an online production environment? Sec. 5.3 provides results to answer RQ1. Sec. 5.4 provides an ablation study to answer RQ2. Sec. 5.6 provides online performance results to answer RQ3.
5.1. Experimental Setup
5.1.1. Data Description
Shanghai Qingdao
[…]
PQ IQ Match
[…]
PQ IQ Match
IP (Gunawan et al., 2014) 6.4 1573.3 2.96 / 30.3 26.2 17.8 7.6 4134.3 2.86 / 23.6 16.8 20.2
Ernie-Bot 4.0 (Sun et al., 2021) 15.7 513.5 0.91 15.2 42.1 46.5 42.5 27.2 776.2 0.78 21.4 43.4 38.2 33.3
GPT-3.5 (OpenAI, [n. d.]) 16.6 422.4 0.83 13.5 40.4 43.1 45.4 25.5 691.5 0.55 22.0 33.4 39.0 46.6
GPT-4 (OpenAI, 2023) 18.0 267.2 0.56 8.2 45.0 48.2 46.9 27.3 569.4 0.49 19.6 46.6 48.7 48.4
GPT-4 CoT (Wei et al., 2022) 18.4 258.3 0.49 7.5 / / / 30.2 542.6 0.43 17.8 / / /
Ours 31.4 86.0 0.42 / 69.8 64.6 72.0 35.4 225.8 0.26 / 71.2 68.2 67.8
Beijing Hangzhou
[…]
PQ IQ Match
[…]
PQ IQ Match
IP (Gunawan et al., 2014) 3.3 3034.2 2.26 / 27.8 18.2 20.4 1.8 1744.4 1.52 / 34.8 31.4 22.5
Ernie-Bot 4.0 (Sun et al., 2021) 18.8 379.4 0.74 12.8 31.2 34.8 32.1 12.9 605.2 1.17 24.4 43.6 34.3 38.2
GPT-3.5 (OpenAI, [n. d.]) 19.7 347.8 0.58 14.3 29.2 40.5 43.8 14.4 665.4 1.16 19.8 41.2 40.8 32.8
GPT-4 (OpenAI, 2023) 20.6 342.6 0.52 11.1 45.4 43.6 45.2 14.8 746.1 1.09 23.2 46.2 39.6 39.4
GPT-4 CoT (Wei et al., 2022) 21.0 327.7 0.54 10.2 / / / 15.5 455.0 1.09 18.6 / / /
Ours 28.4 79.2 0.46 / 59.2 67.6 75.2 21.4 30.5 0.12 / 61.6 65.4 68.3
Table 1. Overall results. GPT-based metrics represent the win rate against GPT-4 CoT.

### v4

@@ 3.4 Preference-aware POI Retrieval
Ideally, we want the queried POIs to best fit the positive subrequest while avoiding the negative subrequest. To achieve this, we use the positive embedding $\mathbf{e}^{\operatorname{pos}}_{i}$ to retrieve $k$ POIs from $\mathcal{P}$ to obtain $\mathcal{P}^{\operatorname{pos}}_{i}=\{p^{\operatorname{pos}}_{i,j}\}_{j=1}^{k}$ and the corresponding embeddings $\mathcal{E}_{i}^{\operatorname{pos}}=\{\mathbf{e}^{\operatorname{pos}}_{i,j}\}_{j=1}^{k}$ from $\mathcal{E}$ with top similarity scores $\mathcal{S}^{\operatorname{pos}}_{i}=\{s^{\operatorname{pos}}_{i,j}\}_{j=1}^{k}$, where $p^{\operatorname{pos}}_{i,j}$ and $s^{\operatorname{pos}}_{i,j}$ represent the $j$th POI and score for $i$th positive sub-request. Next, we compute the similarity scores between $\mathcal{E}^{\operatorname{pos}}_{i}$ and $\mathbf{e}^{\operatorname{neg}}_{i}$ and rerank the POIs based on the gap between positive and negative similarity scores. Using $\mathcal{E}\in\mathbb{R}^{N\times d}$ to denote the pre-computed POI embeddings in the user-owned database, the process is:
$\mathcal{P}^{\operatorname{pos}}_{i},\mathcal{S}^{\operatorname{pos}}_{i},\mathcal{E}^{\operatorname{pos}}_{i}$ $=$ $\operatorname{score}^{k}\left(\mathbf{e}^{\operatorname{pos}}_{i},\mathcal{E}\right)$ (1)
$\mathcal{P}^{\operatorname{neg}}_{i},\mathcal{S}^{\operatorname{neg}}_{i},\mathcal{E}^{\operatorname{neg}}_{i}$ $=$ $\operatorname{score}\left(\mathbf{e}^{\operatorname{neg}}_{i},\mathcal{E}^{\operatorname{pos}}_{i}\right)$ (2)
$\mathcal{P}_{i},\mathcal{S}_{i}$ $=$ $\operatorname{rank}\left(\mathcal{P}^{\operatorname{pos}}_{i},\mathcal{S}^{\operatorname{pos}}_{i}-\mathcal{S}^{\operatorname{neg}}_{i}\right),$ (3)
@@
where the $\operatorname{score}(\cdot)$ function measures embedding similarities, and the superscript $k$ indicates it returns the top-$k$ results . The $\operatorname{rank}(\cdot)$ reorders POIs from highest to lowest similarity scores.
@@
Lastly, we select the top-$k$ POIs with the highest summed scores from the union of all retrieved POIs to form the final set $\mathcal{P}^{rt}$ for the user request $r$:
$\mathcal{P}^{rt},\mathcal{S}^{rt}=\operatorname{rank}^{k}\left(\cup_{i=1}^{|\mathcal{R}|}\left(\mathcal{P}_{i},\mathcal{S}_{i}\right)\right)\cup\left(\mathcal{P}^{\operatorname{must}},\mathcal{S}^{\operatorname{must}}\right),$ (4)
@@
where $\mathcal{S}^{\operatorname{must}}$ has large values to ensure must-see POIs are included. During the union process, scores for the same POI under different subrequests are summed to obtain the final score.
Shanghai Qingdao
Method Rule-based Metrics LLM-Eval $\uparrow$ (%) Rule-based Metrics LLM-Eval $\uparrow$ (%)
$R$ $\uparrow$
AM $\downarrow$
OL $\downarrow$
FR $\downarrow$
PQ IQ Match
$R$ $\uparrow$
AM $\downarrow$
OL $\downarrow$
FR $\downarrow$
PQ IQ Match
Ground Truth / 124.4 0.44 / 68.9 61.5 80.9 / 356.6 0.31 / 75.4 63.9 71.4
IP 6.4 1573.3 2.96 / 30.3 26.2 17.8 7.6 4134.3 2.86 / 23.6 16.8 20.2
Ernie-Bot 4.0 15.7 513.5 0.91 15.2 42.1 46.5 42.5 27.2 776.2 0.78 21.4 43.4 38.2 33.3
GPT-3.5 16.6 422.4 0.83 13.5 40.4 43.1 45.4 25.5 691.5 0.55 22.0 33.4 39.0 46.6
GPT-4 18.0 267.2 0.56 8.2 45.0 48.2 46.9 27.3 569.4 0.49 19.6 46.6 48.7 48.4
GPT-4 CoT 18.4 258.3 0.49 7.5 / / / 30.2 542.6 0.43 17.8 / / /
ItiNera (ours) 31.4 86.0 0.42 / 69.8 64.6 72.0 35.4 225.8 0.26 / 71.2 68.2 67.8
Beijing Hangzhou
Method Rule-based Metrics LLM-Eval $\uparrow$ (%) Rule-based Metrics LLM-Eval $\uparrow$ (%)
$R$ $\uparrow$
AM $\downarrow$
OL $\downarrow$
FR $\downarrow$
PQ IQ Match
$R$ $\uparrow$
AM $\downarrow$
OL $\downarrow$
FR $\downarrow$
PQ IQ Match
Ground Truth / 218.3 0.53 / 61.9 57.3 77.0 / 70.9 0.34 / 47.5 53.2 66.3
IP 3.3 3034.2 2.26 / 27.8 18.2 20.4 1.8 1744.4 1.52 / 34.8 31.4 22.5
Ernie-Bot 4.0 18.8 379.4 0.74 12.8 31.2 34.8 32.1 12.9 605.2 1.17 24.4 43.6 34.3 38.2
GPT-3.5 19.7 347.8 0.58 14.3 29.2 40.5 43.8 14.4 665.4 1.16 19.8 41.2 40.8 32.8
GPT-4 20.6 342.6 0.52 11.1 45.4 43.6 45.2 14.8 746.1 1.09 23.2 46.2 39.6 39.4
GPT-4 CoT 21.0 327.7 0.54 10.2 / / / 15.5 455.0 1.09 18.6 / / /
ItiNera (ours) 28.4 79.2 0.46 / 59.2 67.6 75.2 21.4 30.5 0.12 / 61.6 65.4 68.3
Table 1: Overall results on four datasets. LLM-evaluated metrics are win rates against GPT-4 CoT.
3.5 Cluster-aware Spatial Optimization
3.5.1 Spatial Clustering and Filtering
Algorithm 1 Spatial Clustering & Candidate POI Selection
@@ 3.5.1 Spatial Clustering and Filtering
1:Input: Retrieved POI set $\mathcal{P}^{rt}$ with scores $\mathcal{S}^{rt}$, Distance threshold $\tau$, Candidate POIs num threshold $N^{c}$
2:Output: Spatial Clusters $\mathcal{C}$, Candidate POIs $\mathcal{P}^{c}$
@@
5:for $p_{a}^{rt},p_{b}^{rt}\in V$ with $a\neq b$ do
6:  if ${distance}(p_{a}^{rt},p_{b}^{rt})<\tau$ then
7:   $E\leftarrow E\cup\{(p_{a}^{rt},p_{b}^{rt})\}$
@@
10:while $V\neq\emptyset$ do
11:  $c\leftarrow\text{largest clique in }G$
12:  $\mathcal{C}\leftarrow\mathcal{C}\cup\{c\}$; $V\leftarrow V\setminus c$
@@
15:for each cluster $c\in\mathcal{C}$ do
16:  $s^{rt}_{c}\leftarrow\sum_{p_{j}\in c}s_{j}^{rt}$
@@
18:Sort $\mathcal{C}$ in descending order of $\mathcal{S}^{c}=\{s^{rt}_{c}\}_{c=1}^{\mathcal{C}}$
19:while $|\mathcal{P}^{c}|<N^{c}$ do
20:  $c_{\text{max}}\leftarrow\arg\max_{c\in\mathcal{C}}s_{c}^{rt}$
21:  $\mathcal{P}^{c}\leftarrow\mathcal{P}^{c}\cup c_{\text{max}}$; $\mathcal{C}\leftarrow\mathcal{C}\setminus\{c_{\text{max}}\}$
@@
23:return $\mathcal{C},\mathcal{P}^{c}$
A spatially coherent itinerary enhances the travel experience by allowing travelers to move efficiently between clusters of POIs, reducing transit time and effort Bolzoni et al. (2014). Therefore, spatially filtering and sequencing the retrieved POIs is essential. To achieve this, we compute spatial clusters of the retrieved POIs and select candidates based on proximity and matching scores, addressing cluster-aware spatial optimization by solving a hierarchical traveling salesman problem Jiang et al. (2014). Given the retrieved POIs $\mathcal{P}^{rt}$, we create a spatial proximity graph $G$ using a distance matrix $D$, where each node is a POI and edges connect nodes within a distance threshold $\tau$. A community detection algorithm divides $G$ into clusters. We iteratively select the cluster with the highest summed similarity score until the total number of selected POIs reaches a threshold $N^{c}$, forming the candidate POIs $\mathcal{P}^{c}$ for the user request $r$. The process is detailed in Algo. 1.
3.5.2 POI Ordering via Solving Hierarchical TSP
@@
1:Input: Spatial clusters $\mathcal{C}$, candidate POIs $\mathcal{P}^{c}$, distance matrix $D$
2:Output: Ordered list of candidate POIs $\mathcal{P}^{\text{order}}$
@@
5:for each cluster $c$ in $\mathcal{C}^{\text{order}}$ do
6:  $p_{\text{start}}^{c},p_{\text{end}}^{c}\leftarrow\text{GetClusterEndpoints}(c,\mathcal{C}^{\text{order}},D)$
7:  $c_{\text{path}}\leftarrow\text{SolveConstraintTSP}(c,p_{\text{start}}^{c},p_{\text{end}}^{c},D)$
8:  $\mathcal{P}^{\text{order}}\leftarrow\mathcal{P}^{\text{order}}\cup c_{\text{path}}$
@@
13:return $\mathcal{P}^{\text{order}}$
After obtaining the spatial clusters $\mathcal{C}$, we optimize the access order of the candidate POIs for a spatially coherent itinerary by determining the access order of each cluster and solving TSPs within each cluster with start and end POI constraints. Start and end points are selected based on the proximity of POIs in adjacent clusters, as shown in Fig. 2. This process, outlined in Algo. 2, optimizes and ensures coherent traversal among selected POIs using an efficient hierarchical TSP approach. ‘SolveTSP’ and ‘SolveTSPWithEndpoints’ handle standard and constrained TSPs, respectively, while ‘GetClusterEndpoints’ determines the start and end points for each cluster. The starting point, $p_{\text{start}}$, is identified by ‘Select’ from subrequests $\mathcal{R}$ or by prompting an LLM with $\mathcal{P}^{c}$ and $r$. Finally, the ‘Reorder’ function arranges the POIs in the original order of precedence starting from $p_{\text{start}}$. Further details are in §D.
3.6 Itinerary Generation
@@
Selecting a subset from $\mathcal{P}^{\text{order}}$ ensures a spatially coherent itinerary, but a high-quality itinerary must also meet constraints like time availability and practicality. It should, for example, avoid consecutive restaurant visits and schedule activities appropriately, such as bars in the evening or coffee shops in the morning. Traditional optimization algorithms can become overly complex and lack variability Yochum et al. (2020); Taylor et al. (2018), hindering itinerary diversity. To address this, we leverage the advanced reasoning and planning capabilities of LLMs to generate final itineraries that meet these diverse constraints.
@@
Specifically, the primary objective of this module is to effectively utilize LLM to select an optimal subset from $\mathcal{P}^{\text{order}}$, which closely aligns with user requests while adhering to various constraints. This process can be formally defined as follows:
$I\sim\operatorname{LLM}\left(\mathbb{P}_{IG}\left(\mathbf{r},\,\mathcal{P}^{\text{order}}\,,I_{ex}\right)\right),$ (5)
where $I_{ex}$ indicates extra natural language input that improves the language quality of the generated itinerary. The prompt $\mathbb{P}_{IG}$ for generating the final itinerary contains the following parts: (1) User request information; (2) Candidate POIs with context; (3) Task description; (4) Specific constraints; (5) Language style; (6) Output format. The full prompt is provided in §F.4.
4 Experiments
4.1 Experimental Setup